OpenStudy (anonymous):

use l'hopital's rule to find the limit. lim ln(x^2 + 11x) / lnx x->0

OpenStudy (anonymous):

differentiate top and bottom

OpenStudy (anonymous):

= limit [ (2x+11) / (x^2 +11x) ] / [1/x]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}= \frac{x (2x +11) }{x^2 +11x} \]

OpenStudy (anonymous):

\[= \lim_{x \rightarrow 0} \frac{2x^2 +11x}{x^2 +11x}\]

OpenStudy (anonymous):

still 0/0 differentiate again

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}= \frac{4x + 11}{2x+11} \]

OpenStudy (anonymous):

now, sub in x=0 you get 11/11 = 1 limit is 1